# LaDTalk: Latent Denoising for Synthesizing Talking Head Videos with High Frequency Details

### v2

@@ 5.3 Ablation Study
VQ Regularization Loss. As previously discussed, in practical applications, training without regularization on the VQ codebook can result in a decrease in noise robustness due to the stochastic optimization of the codebook. To illustrate the efficacy of  Eq. 11 in stabilizing this robustness, we conducted a comparison between the VQAE with and without codebook optimization. As shown in  Fig. 8, the absence of our regularization loss introduces numerous artifacts in the final video, attributable to errors in latent matching.
To further demonstrate the necessity and efficacy of our VQ regularization loss, we conduct detailed ablation studies on the lower bound and optimization settings. As shown in Tab. 3, optimizing the minimal distance pair in the codebook yields the best results: (1) Optimizing pairwise distances within the codebook leads to codebook collapse, as indicated by the first two rows. (2) In contrast, local optimization with $\theta=1$ is effective. (3) Increasing the distance ($\theta=2$) in local optimization degrades performance due to over-constraint. This contrasts with general principles for training VQ-VAEs, where maintaining a sufficiently large distance in the codebook is crucial for diversity in image representation. However, our goal differs: while VQ-VAE aims for diverse codebooks to handle varied images, SOVQAE focuses on enhancing noise robustness for identity-specific faces. We do not explore $\theta<1$ since $d_{c}<1$ is typically the case.
@@
Length of training videos. An essential inquiry pertains to the optimal length of video required to train a sufficiently expressive SOVQAE to support our method in achieving realistic and temporally consistent denoising performance. To address this question, we experimentally varied the length of the training video for the SOVQAE and assessed the corresponding impact on performance. The findings, as detailed in  Tab. 4, yield the following insights: Using extended training videos improves the audio-driven capabilities of SOVQAE. A 5-minute training video is adequate for generating realistic talking head videos with our method, indicating similar data requirements to NeRF-based methods. To make our manuscript more sufficient, we introduce our limitations in Appendix B.
@@ 6 Conclusion
In this paper, we introduce LaDTalk, a novel method for ID-specific talking head generation that achieves state-of-the-art performance. By leveraging the theory of Lipschitz Continuity, we have theoretically established noise robustness in VQAEs. Building on this foundation, we propose SOVQAE, a plug-and-play denoising model designed to achieve temporally consistent recovery of high-frequency details from low-quality outputs, such as those generated by Wav2Lip. Through extensive experimental validation, we demonstrate that LaDTalk attains new SOTA video quality and out-of-domain lip synchronization accuracy. Additionally, we showcase the versatility of SOVQAE across various base models. This work highlights the potential of post-processing paradigms in the domain of talking head generation. We hope that our contributions will provide valuable insights to the research community.
Supplementary Material
@@
In our experiments, we configured the codebook with $c=256$ channels and a size of $N=1024$. For the Usual Dataset, we trained the optimized VQAE on face crops resized to a resolution of $256\times 256$ pixels, completing 50 epochs using an NVIDIA 4090 GPU, which amounted to 8.5 GPU hours. For the HFTK Dataset, the face crops were resized to a higher resolution of $512\times 512$ pixels, and the model was trained for 50 epochs on the same hardware, consuming 34 GPU hours. We utilized a downsampling factor of $h_{o}=h/16$ and $w_{o}=w/16$. Additional details regarding the network architecture are presented in our Appendix F. The foundational Wav2Lip model employed in our experiments is a pretrained version sourced from the official repository [^2] https://github.com/Rudrabha/Wav2Lip. For the regularization loss defined in  Eq. 11, we have empirically set $\theta=1$. The high frequency videos consists of 4 4K videos from YouTube, reprocessed to $1024\times 1024$ at 25 fps with a 10M bitrate (compared to 3M in the Usual dataset). They have higher resolution and richer high-frequency textures, as shown in Fig. 9. These videos enable evaluation of identity-specific methods under more detailed and realistic conditions, aligning with industrial needs.
@@ Appendix B Discussion
Limitations. In comparison with existing methods, our approach has a notable limitation in terms of computational requirements for both training and inference phases. While NeRF-based methods such as SyncTalk [2] and GeneFace++ can achieve real-time inference, our method operates at a 1:1.2 inference ratio on an Nvidia 4090 GPU when processing videos at 25 FPS, which indicates that it does not yet meet the criteria for real-time performance. Additionally, our training demands are more substantial; for instance, to train our model on 5 minutes of talking head video data, we require approximately 8.5 GPU hours using an NVIDIA 4090 GPU. These computational demands highlight areas for potential future optimization.
@@
Boarder Impact. In contrast to traditional end-to-end frameworks, our method generates realistic talking head videos through a temporally-consistent post-processing approach. From an alternative viewpoint, we also highlight the potential of harnessing the robust lip synchronization capabilities of a pretrained Wav2Lip model. This suggests that integrating a pretrained Wav2Lip model into a NeRF-based talking head pipeline could significantly enhance per-frame stability and synchronization performance. Moreover, a primary contribution of our work is the theoretical proof and practical demonstration of the noise robustness inherent in the Vector Quantization mechanism. Consequently, it is a logical next step to explore the application of this concept within the context of adversarial attacks [48, 49, 50] to bolster the security and reliability of neural networks.
@@
Future works. To the limitation on computational demand, we consider to explore more efficient network design, which guarantees the noise robustness while decrease the size of whole pipeline, aiming to real-time inference and low training burden. Except for the pipeline optimization, we plan to explore the way of prior integration in talking head task. Specifically, the LQ talking head results of Wav2Lip model are seem as a kind of intermediate representation (like the face keypoints representation in GeneFace and GeneFace++). Following these works’ idea, integrating such more semantic-rich intermediate representation into NeRF or 3DGS pipeline seems like a promising way to achieve better audio generality. Besides, we also notice that SOVQAE incurs decrease of lip-synchronization. Hence, alleviating this phenomena via more delicate network design is also our goal.
@@ Appendix C Additional Experiments
To demonstrate the plug-and-play performance of our framework, we integrated SOVQAE into various base models. As shown in Tab. 5, SOVQAE improves visual quality metrics across different ID-specific base models, confirming its versatility and effectiveness. Notably, LaDTalk achieves the best overall performance by leveraging the exceptional audio-lip synchronization capabilities of Wav2Lip. This superior synchronization positively impacts PSNR and FID values, thereby enhancing overall video quality and ensuring LaDTalk’s top performance.
@@
Table 5: Additional experiments in Usual Dataset.
@@
GeneFace++ 31.11 20.55 6.89 7.50
GeneFace++ + SOVQAE 33.27 18.21 6.42 8.22
SyncTalk (CVPR 2024) 36.06 6.49 7.05 7.68
SyncTalk + SOVQAE 37.16 5.12 6.58 7.88
Ours 38.97 3.45 8.13 6.44
@@ D.2 One layer CNN
We refer to [28, Part 3] for details.
@@
where $\rho(A)$ is the spectral radius of $A$. See details for [29, Theorem 6.15].
@@
We refer to [27, Lemma 4.1] for details.
@@ D.3 Multiple layers CNN
Now we mainly consider a CNN as a Lipschitzian continuous function $\epsilon$ with Lipschitzian constant $L_{\epsilon}$. The spaces $\mathcal{S}$ of graphs of size $h\times w$ is a subset of the space $\mathbb{R}^{c_{i}\times h\times w}$. The range $\epsilon(\mathcal{S})$ can be seen as an encoding of $\mathcal{S}$. In this paper, we utilize VQAE to learn suitable encoding of space $\mathcal{S}$ and simultaneously select anchors in this encoding. Our goal is for this network to learn the features of high-quality images and minimize the distance between codebook and the features. Meanwhile, we also wish the network to be robust with mild input disturbance .Next, we will demonstrate the feasibility of our approach.
@@
Let $\{\epsilon,\delta,{\bf C}\subset\mathbb{R}^{c},g_{{\bf C}}\}$ be VQAE with single channel latent space, where $\epsilon$ is the CNN encoder with Lipschitzian constant $L_{\epsilon}$, and $\mathbf{C}\subset\epsilon(\mathcal{S})$ is a set of codebook anchors. Define $d_{\mathbf{C}}=\min\{\|a-b\|_{F}:a,b\in\mathbf{C},a\neq b\}$ and $\gamma$ be the maximal distance of high-quality image latent to closest anchor, that is, $\gamma=\max_{p\in\epsilon(\mathbf{HQ})}d(p,\mathbf{C})$, where $\mathbf{HQ}$ is the space of high quality images, and $p\in\mathbb{R}^{c}$ is single channel vector. Assume that $2\gamma<d_{\mathbf{C}}$ (for almost all cases, it is naturally satisfied). For all low-quality image $y$ corresponding to a high-quality image $x$ such that
@@
If the VQAE $\{\epsilon,\delta,{\bf C}\subset\mathbb{R}^{c},g_{{\bf C}}\}$ is convergent, the distance of high-quality images to codebook $\mathbf{C}$ is extremely small, which means that $\gamma\ll 1$ is negligible.
@@ Appendix F More implementation details
Table 6: High-level architecture of the encoder and decoder of our SOVQAE. The design of the networks follows the architecture presented in [58] with no skip-connections. For the discriminator, we use a patch-based model. Note that $h=\frac{H}{2^{m}},w=\frac{W}{2^{m}}$ and $f=2^{m}$.

### v3

@@ 5.3 Ablation Study
VQ Regularization Loss. As previously discussed, in practical applications, training without regularization on the VQ codebook can result in a decrease in noise robustness due to the stochastic optimization of the codebook. While Fig. 1 demonstrates impressive robustness under Gaussian blur and Gaussian noise, vanilla VQ-AE falls short to handle the unknown noise disturbance between the output of Wav2Lip and ground truth, resulting artifacts. To illustrate the efficacy of Eq. 7 in stabilizing this robustness, we conducted a comparison between the VQ-AE with and without codebook optimization. As shown in Fig. 7, the absence of our regularization loss introduces numerous artifacts in the output video, attributable to errors in latent matching.
@@
To further demonstrate the necessity and efficacy of our VQ regularization loss, we conduct detailed ablation studies on the lower bound $\theta$ and optimization settings. As shown in Tab. 3, optimizing the minimal distance pair in the codebook yields the best results: (1) Optimizing pairwise distances within the codebook leads to codebook collapse, as indicated by the first two rows. (2) In contrast, local optimization with $\theta=1$ is effective. (3) Increasing the distance ($\theta=2$) in local optimization degrades performance due to over-constraint. This contrasts with general principles for training VQ-VAEs, where maintaining a sufficiently large distance in the codebook is crucial for diversity in image representation. However, our goal differs: while VQ-VAE aims for diverse codebooks to handle varied images, SOVQAE focuses on enhancing noise robustness for identity-specific faces. We do not explore $\theta<1$ since $d_{c}<1$ is typically the case.
@@
Length of training videos. An essential inquiry pertains to the least length of video required to train a sufficiently expressive SOVQAE to support our pipeline in achieving realistic and temporally consistent denoising performance. To address this question, we experimentally varied the length of the training video for the SOVQAE and assessed the corresponding impact on performance. The findings, as detailed in Tab. 4, yield the following insights: Using extended training videos improves the audio-driven capabilities of SOVQAE. A 5-minute training video is adequate for generating realistic talking head videos with our method, indicating similar data requirements to NeRF-based methods. To make our manuscript more sufficient, we introduce our limitations in Appendix B.
@@ 6 Conclusion
This work first establishes a theoretical foundation for the noise robustness of VQ-AE by proving an upper bound based on Lipschitz continuity theory, which enables subsequent method design. Building on this foundation, we propose SOVQAE, enhancing temporal consistency in latent space via a space regularization loss. Furthermore, we introduce an efficient cascade pipeline combining a pre-trained Wav2Lip model with ID-specific SOVQAE, achieving state-of-the-art video quality and out-of-domain lip synchronization accuracy. The whole pipeline requires only 5 GPU hours (NVIDIA 4090) for training and supports 30 FPS real-time inference, making it suitable for industrial deployment. In other words, this work highlights the potential of post-processing paradigms in the domain of talking head generation. We hope that our contributions will provide valuable insights to the research community.
@@
In our experiments, we configured the codebook with $c=256$ channels and a size of $N=1024$. For the Usual Dataset, we trained the optimized VQ-AE on face crops resized to a resolution of $256\times 256$ pixels, completing 50 epochs using an NVIDIA 4090 GPU, which amounted to 5 GPU hours. For the HF videos, the face crops were resized to a higher resolution of $512\times 512$ pixels, and the model was trained for 50 epochs on the same hardware, consuming 20 GPU hours. We utilized a downsampling factor of $h_{o}=h/16$ and $w_{o}=w/16$. Additional details regarding the network architecture are presented in our Appendix F. The foundational Wav2Lip model employed in our experiments is a pretrained version sourced from the official repository [^2] https://github.com/Rudrabha/Wav2Lip. For the regularization loss defined in Eq. 7, we have empirically set $\theta=1$. The high frequency videos consists of 4 4K videos from YouTube, reprocessed to $1024\times 1024$ at 25 fps with a 10M bitrate (compared to 2M in the Usual dataset). They have higher resolution and richer high-frequency textures, as shown in Fig.8. These videos enable evaluation of identity-specific methods under more detailed and realistic conditions, aligning with industrial needs.
@@
Figure 9: Illustration of our Sliding Evaluation method.
Sliding Evaluation. In our experiments, we observe that the generated video lengths from various methods prajwal2020lip ; zhang2023dinet ; ye2023geneface ; peng2023synctalk ; li2023efficient are not only disparate but also consistently shorter than those of the ground truth videos. Directly assessing image quality in conjunction with video duration could potentially diminish the perceived performance of these methods. To address this issue, we introduce a Sliding Evaluation technique. As illustrated in Fig. 9, this approach treats the generated video sequence as a dynamic sliding window, calculating the average image quality across all frames captured within the window. The window progressively moves along the timeline, and the maximum value obtained from these iterations is adopted as the final evaluation outcome. We apply this Sliding Evaluation method to all PSNR and FID results to evaluate different methods, thus ensuring a more holistic and equitable assessment.
@@ B.1 Limitations
In comparison with existing methods, our approach has some limitations in terms of computational requirements for both training and inference phases. While NeRF-based methods such as SyncTalk peng2023synctalk and GeneFace++ can achieve faster inference (>40 FPS), our pipeline has only 30 FPS inference speed on an Nvidia 4090 GPU. Besides, some 3DGS-based methods could even achieve ultra inference speed (>100 FPS). Additionally, our training demand is more substantial; for instance, to train our model on 5 minutes of talking head video data, we require approximately 5 GPU hours using an NVIDIA 4090 GPU. These computational demands highlight areas for potential future optimization. On the other hand, our pipeline cannot achieve novel view synthesis, as it is a video post-processing pipeline.
Nevertheless, we believe our approach remains valuable for repairing high-frequency details in identity-specific talking-head generation (ID-specific THG).
B.2 Boarder Impacts
B.2.1 Postive impacts
@@
In contrast to traditional end-to-end frameworks, our method generates realistic talking head videos through a temporally-consistent post-processing approach. From an alternative viewpoint, we also highlight the potential of harnessing the robust lip synchronization capabilities of a pretrained Wav2Lip model. This suggests that integrating a pretrained Wav2Lip model into a NeRF-based talking head pipeline could significantly enhance per-frame stability and synchronization performance. Moreover, a primary contribution of our work is the theoretical proof and practical demonstration of the noise robustness inherent in the Vector Quantization mechanism. Consequently, it is a logical next step to explore the application of this concept within the context of adversarial attacks madry2017towards ; xu2020adversarial ; zeng2019adversarial to bolster the security and reliability of neural networks.
B.2.2 Negative impacts.
Our work proposes a method for generating high-fidelity, identity-specific talking-head videos. While such technology has promising applications in entertainment, virtual assistants, and telecommunication, it also raises critical ethical concerns. For example, our method could be misused to create highly realistic fake videos for malicious purposes, such as identity theft, political manipulation, or financial fraud. The realism of ID-specific generation may reduce public trust in video content. Hence, if our model is released publicly, we will require users to agree to a usage policy prohibiting misuse (e.g., deepfake generation, identity fraud). Access may be restricted to verified researchers or institutions.
B.2.3 Future works.
@@
To the limitation on computational demand, we consider to explore more efficient network design, which guarantees the noise robustness while decrease the size of whole pipeline, aiming to faster inference and lower training burden. Except for the pipeline optimization, we plan to explore the way of prior integration in talking head task. Specifically, the LQ talking head results of Wav2Lip model are seem as a kind of intermediate representation (like the face keypoints representation in GeneFace and GeneFace++). Following these works’ idea, integrating such more semantic-rich intermediate representation into NeRF or 3DGS pipeline seems like a promising way to achieve better audio generality. Besides, we also notice that SOVQAE incurs decrease of lip-synchronization. Hence, alleviating this phenomena via more delicate network design is also our goal. Besides, it is notable that SOVQAE still has mild damage in lip-synchronization metric to the output of Wav2Lip model, maybe we can explore an extra lightweight network to improve this.
@@ Appendix C Additional Experiments
C.1 Plug-and-Play Experiment
To demonstrate the plug-and-play performance of our SOVQAE, we integrated SOVQAE into various base models. In this experiment, we follow our cascade pipeline workflow: 1) conduct face detection from output video from different base models; 2) resize the detected faces to $256\times 256$ then put them into SOVQAE; 3) resize back to detected shape and paste back to output video from base model. As shown in Tab.2, SOVQAE improves visual quality metrics across different ID-specific base models, confirming its versatility and effectiveness. Notably, our cascade pipeline achieves the best overall performance by leveraging the exceptional audio-lip synchronization capabilities of Wav2Lip. This superior synchronization positively impacts PSNR and FID values, thereby enhancing overall video quality and ensuring cascade pipeline’s top performance.
@@ C.1 Plug-and-Play Experiment
C.2 OOD Audio-driven Experiment
As shown in Table 5, we present a detailed breakdown of each item in our out-of-distribution (OOD) audio-driven experiment. The evaluation covers 10 diverse audio clips spanning multiple languages—including Japanese, English, Chinese, Spanish, and French—as well as varying speaker ages and genders. Across all test cases, our pipeline consistently achieves the best lip synchronization performance.
Method DINetzhang2023dinet IP-LAPzhong2023identity GeneFace++ye2023geneface++ SyncTalkpeng2023synctalk Ours
@@ C.2 OOD Audio-driven Experiment
Audio 1 4.7768 8.9890 4.4024 9.3161 4.8974 8.9521 4.2177 9.28129 6.2354 8.0987
Audio 2 4.3692 9.0287 4.1531 8.6423 4.6195 8.7908 4.6328 8.0043 5.9480 7.6968
Audio 3 5.6343 8.6510 4.8950 9.1588 5.1553 8.8068 6.3866 8.0077 7.4764 7.5445
Audio 4 4.2127 8.5808 4.8710 8.6493 3.9984 8.2326 5.7213 7.8372 6.7817 7.5449
Audio 5 6.1828 8.5279 3.8994 9.4514 5.9873 8.5823 4.4513 8.8178 6.0246 7.9151
Audio 6 6.2382 7.9339 5.1737 8.5354 5.9934 7.7590 6.0401 7.1978 7.1988 7.3960
Audio 7 4.6975 8.7730 4.6781 9.2968 5.0774 8.3832 5.4807 8.2488 7.2357 7.1888
Audio 8 5.5323 8.3505 4.5502 8.9811 5.9367 7.8090 5.4721 7.9490 6.8894 7.2750
Audio 9 5.0672 8.8984 4.7241 9.2790 5.5664 8.6923 5.7529 7.7905 6.9524 7.5649
Audio 10 5.5989 8.1889 2.4123 9.2861 6.2121 7.9188 2.9075 7.3401 3.7788 8.4436
Best.
@@ D.2 One layer CNN
We refer to (dummit2004abstract, , Part 3) for details.
@@
where $\rho(A)$ is the spectral radius of $A$. See details for (shores2007applied, , Theorem 6.15).
@@
We refer to (gray2006toeplitz, , Lemma 4.1) for details.
@@ D.3 Multiple layers CNN
Now we mainly consider a CNN as a Lipschitzian continuous function $\epsilon$ with Lipschitzian constant $L_{\epsilon}$. The spaces $\mathcal{S}$ of graphs of size $h\times w$ is a subset of the space $\mathbb{R}^{c_{i}\times h\times w}$. The range $\epsilon(\mathcal{S})$ can be seen as an encoding of $\mathcal{S}$. In this paper, we utilize VQ-AE to learn suitable encoding of space $\mathcal{S}$ and simultaneously select anchors in this encoding. Our goal is for this network to learn the features of high-quality images and minimize the distance between codebook and the features. Meanwhile, we also wish the network to be robust with mild input disturbance .Next, we will demonstrate the feasibility of our approach.
@@
Let $\{\epsilon,\delta,{\bf C}\subset\mathbb{R}^{c},g_{{\bf C}}\}$ be VQ-AE with single channel latent space, where $\epsilon$ is the CNN encoder with Lipschitzian constant $L_{\epsilon}$, and $\mathbf{C}\subset\epsilon(\mathcal{S})$ is a set of codebook anchors. Define $d_{\mathbf{C}}=\min\{\|a-b\|_{F}:a,b\in\mathbf{C},a\neq b\}$ and $\gamma$ be the maximal distance of high-quality image latent to closest anchor, that is, $\gamma=\max_{p\in\epsilon(\mathbf{HQ})}d(p,\mathbf{C})$, where $\mathbf{HQ}$ is the space of high quality images, and $p\in\mathbb{R}^{c}$ is single channel vector. Assume that $2\gamma<d_{\mathbf{C}}$ (for almost all cases, it is naturally satisfied). For all low-quality image $y$ corresponding to a high-quality image $x$ such that
@@
If the VQ-AE $\{\epsilon,\delta,{\bf C}\subset\mathbb{R}^{c},g_{{\bf C}}\}$ is convergent, the distance of high-quality images to codebook $\mathbf{C}$ is extremely small, which means that $\gamma\ll 1$ is negligible.
@@ Appendix F More implementation details
Table 6: High-level architecture of the encoder and decoder of our SOVQAE. The design of the networks follows the architecture presented in isola2017image with no skip-connections. For the discriminator, we use a patch-based model. Note that $h=\frac{H}{2^{m}},w=\frac{W}{2^{m}}$ and $f=2^{m}$.